OpenStudy (anonymous):

Find the range of the exponential function f(x) = -2(3)x

OpenStudy (anonymous):

@ranga Can you help have couple of questions

OpenStudy (ranga):

I can try.

OpenStudy (ranga):

Is that function f(x) = -6x ?

OpenStudy (anonymous):

Its asking if its positive real numbers all real numbers...

OpenStudy (anonymous):

negative and non-zero real numbers

OpenStudy (ranga):

Was the function typed correctly? f(x) = -2(3)x = -6x ??

OpenStudy (anonymous):

ohh no ^x my bad

OpenStudy (anonymous):

f(x)=-2(3)^x

OpenStudy (ranga):

oh, ok. It did not look right. f(x) = -2(3^x). We can try a few values for x and evaluate f(x) x f(x) 0 -2 1 -6 2 -18 infinity -infinity -1 -2/3 -2 -2/9 -3 -2/27 -infinity 0

OpenStudy (anonymous):

negative real numbers?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

Thanks got cople more questioncs and im done can you help or are you buisy?

OpenStudy (ranga):

I have about 10 minutes and I have to leave after that.

OpenStudy (anonymous):

OK